Select the ordered pair from the choices below that is a solution to the following system of equations:
8x - 2 = y
5y - 12x = -3
A. (1, 6)
B. (1/4, 0)
C. (-2, -18)
D. (1/2, 2)
Please help !

Answer by Cromlix(4381)

y - 8x = -2...1
5y - 12x = -3...2
Multiply (1) by 5
5y - 40x = -10...3
5y - 12x = -3....2
Subtract (2) from (3)
-28x = -7
x = 1/4
Substitute x = 1/4 in (1)
y - 8(1/4) = -2
y - 2 = -2
y = 0
{1/4, 0}
B
Hope this helps.
:-)